# Diffusion-based Dynamic Contract for Federated AI Agent Construction in Mobile Metaverses

### v1

@@ Appendix B Proof for Lemma 7
$\begin{split}&u_{k}^{1}(T_{k}^{1})+\beta\sum_{j=2}^{K}p_{j}^{2}(T_{k}^{1})\\ &\qquad\times\Bigg{[}\frac{\theta_{j}^{2}E}{\theta_{1}^{2}}+\sum_{l=1}^{j-1}\theta_{j}^{2}c\bigg{(}\frac{T_{l}^{2}(\theta_{k}^{1})}{\theta_{l}^{2}}-\frac{T_{l}^{2}(\theta_{k}^{1})}{\theta_{l+1}^{2}}\bigg{)}\Bigg{]}\\ &\quad\geq u_{k}^{1}(T^{1}_{k^{\prime}})+\beta\sum_{j=2}^{K}p_{j}^{2}(T_{k}^{1})\\ &\qquad\times\Bigg{[}\frac{\theta_{j}^{2}E}{\theta_{1}^{2}}+\sum_{l=1}^{j-1}\theta_{j}^{2}c\bigg{(}\frac{T_{l}^{2}(\theta^{1}_{k^{\prime}})}{\theta_{l}^{2}}-\frac{T_{l}^{2}(\theta^{1}_{k^{\prime}})}{\theta_{l+1}^{2}}\bigg{)}\Bigg{]}.\end{split}$ (B-1)
@@
$a=\frac{\theta_{j}^{2}E}{\theta_{1}^{2}}+\sum_{l=1}^{j-1}\theta_{j}^{2}c\bigg{(}\frac{T_{l}^{2}(\theta_{k}^{1})}{\theta_{l}^{2}}-\frac{T_{l}^{2}(\theta_{k}^{1})}{\theta_{l+1}^{2}}\bigg{)},$ (B-2)
$b=\frac{\theta_{j}^{2}E}{\theta_{1}^{2}}+\sum_{l=1}^{j-1}\theta_{j}^{2}c\bigg{(}\frac{T_{l}^{2}(\theta^{1}_{k^{\prime}})}{\theta_{l}^{2}}-\frac{T_{l}^{2}(\theta^{1}_{k^{\prime}})}{\theta_{l+1}^{2}}\bigg{)}.$ (B-3)
Repeating the above procedure can also obtain the IIC constraint for type-$k^{\prime}$ ESs like (B-1). Then, by combining the two IIC constraints, we have
$\begin{split}u_{k}^{1}(T_{k}^{1})&-u_{k}^{1}(T^{1}_{k^{\prime}})+u^{1}_{k^{\prime}}(T^{1}_{k^{\prime}})-u^{1}_{k^{\prime}}(T^{1}_{k})\\ &+\beta\sum_{j=2}^{K}[p_{j}^{2}(T_{k}^{1})-p_{j}^{2}(T^{1}_{k^{\prime}})](a-b)\geq 0.\end{split}$ (B-4)
To better analyze whether (B-4) is true, we separate (B-4) into three parts, which are given by
$\begin{split}u_{k}^{1}(T_{k}^{1})-u_{k}^{1}(T^{1}_{k^{\prime}})&+u^{1}_{k^{\prime}}(T^{1}_{k^{\prime}})-u^{1}_{k^{\prime}}(T^{1}_{k})\\ &\quad=(\theta_{k}^{1}-\theta^{1}_{k^{\prime}})(R_{k}^{1}-R^{1}_{k^{\prime}}),\end{split}$ (B-5)
$p_{j}^{2}(T_{k}^{1})-p_{j}^{2}(T^{1}_{k^{\prime}}),$ (B-6)
@@
$a-b=\theta_{j}^{2}c\sum_{l=1}^{j-1}\Bigg{[}\bigg{(}\frac{1}{\theta_{l}^{2}}-\frac{1}{\theta_{l+1}^{2}}\bigg{)}(T_{l}^{2}(\theta_{k}^{1})-T_{l}^{2}(\theta^{1}_{k^{\prime}}))\Bigg{]}.$ (B-7)
Since the types of period 1 and period 2 are positively correlated[22, 16], we have $p_{j}^{2}(T_{k}^{1})>p_{j}^{2}(T^{1}_{k^{\prime}})$, i.e., (B-6) is positive. To make (B-4) hold, (B-5) and (B-7) must be positive. Following Lemma 7, (B-5) will be positive if $T_{k}^{1}>T^{1}_{k^{\prime}}$, and (B-7) will be positive if $T_{j}^{2}(T_{k}^{1})>T_{j}^{2}(T^{1}_{k^{\prime}})$. Thus, the proof of Lemma 7 is completed.
Appendix C Proof for Lemma 8
@@
$u_{k}^{1}(T_{k}^{1})+\beta\sum_{j=1}^{K}p_{j}^{2}(T_{k}^{1})[u_{j}^{2}(T_{j}^{2}(T_{k}^{1}))-u_{j}^{2}(T_{j}^{2}(T^{1}_{k^{\prime}}))]\geq u_{k}^{1}(T^{1}_{k^{\prime}}).$ (C-1)
Since $\beta$ and $p_{j}^{2}(T_{k}^{1})$ are small, (C-1) can be simplified into
$u_{k}^{1}(T_{k}^{1})\geq u_{k}^{1}(T^{1}_{k^{\prime}}),$ (C-2)
@@
$\theta_{k}^{1}(R_{k}^{1}-R^{1}_{k^{\prime}})\geq c(T_{k}^{1}-T^{1}_{k^{\prime}}).$ (C-3)
We first prove the sufficiency of Lemma 8, namely if $T_{k}^{1}>T^{1}_{k^{\prime}}$, then $R_{k}^{1}>R^{1}_{k^{\prime}}$. From (C-3), we can know that $R_{k}^{1}>R^{1}_{k^{\prime}}$ since $T_{k}^{1}>T^{1}_{k^{\prime}}$. Thus, the sufficiency is proved.
@@
$u^{1}_{k^{\prime}}(T^{1}_{k^{\prime}})\geq u^{1}_{k^{\prime}}(T^{1}_{k}),$ (C-4)
@@
$\theta^{1}_{k^{\prime}}(R^{1}_{k^{\prime}}-R_{k}^{1})\geq c(T^{1}_{k^{\prime}}-T^{1}_{k}).$ (C-5)
Given $R^{1}_{k^{\prime}}>R_{k}^{1}$, the left-hand side of (C-5) is negative. Thus, we have $T^{1}_{k}-T^{1}_{k^{\prime}}$. The proof of Lemma 8 is completed.

### v2

@@ Appendix G Proof for Lemma 7
$\begin{split}&u_{k}^{1}(T_{k}^{1})+\beta\sum_{j=2}^{K}p_{j}^{2}(T_{k}^{1})\\ &\qquad\times\Bigg[\frac{\theta_{j}^{2}E}{\theta_{1}^{2}}+\sum_{l=1}^{j-1}\theta_{j}^{2}c\bigg(\frac{T_{l}^{2}(\theta_{k}^{1})}{\theta_{l}^{2}}-\frac{T_{l}^{2}(\theta_{k}^{1})}{\theta_{l+1}^{2}}\bigg)\Bigg]\\ &\quad\geq u_{k}^{1}(T^{1}_{k^{\prime}})+\beta\sum_{j=2}^{K}p_{j}^{2}(T_{k}^{1})\\ &\qquad\times\Bigg[\frac{\theta_{j}^{2}E}{\theta_{1}^{2}}+\sum_{l=1}^{j-1}\theta_{j}^{2}c\bigg(\frac{T_{l}^{2}(\theta^{1}_{k^{\prime}})}{\theta_{l}^{2}}-\frac{T_{l}^{2}(\theta^{1}_{k^{\prime}})}{\theta_{l+1}^{2}}\bigg)\Bigg].\end{split}$ (G-1)
@@
$a=\frac{\theta_{j}^{2}E}{\theta_{1}^{2}}+\sum_{l=1}^{j-1}\theta_{j}^{2}c\bigg(\frac{T_{l}^{2}(\theta_{k}^{1})}{\theta_{l}^{2}}-\frac{T_{l}^{2}(\theta_{k}^{1})}{\theta_{l+1}^{2}}\bigg),$ (G-2)
$b=\frac{\theta_{j}^{2}E}{\theta_{1}^{2}}+\sum_{l=1}^{j-1}\theta_{j}^{2}c\bigg(\frac{T_{l}^{2}(\theta^{1}_{k^{\prime}})}{\theta_{l}^{2}}-\frac{T_{l}^{2}(\theta^{1}_{k^{\prime}})}{\theta_{l+1}^{2}}\bigg).$ (G-3)
Repeating the above procedure can also obtain the IIC constraint for type-$k^{\prime}$ ESs like (G-1). Then, by combining the two IIC constraints, we have
$\begin{split}u_{k}^{1}(T_{k}^{1})&-u_{k}^{1}(T^{1}_{k^{\prime}})+u^{1}_{k^{\prime}}(T^{1}_{k^{\prime}})-u^{1}_{k^{\prime}}(T^{1}_{k})\\ &+\beta\sum_{j=2}^{K}[p_{j}^{2}(T_{k}^{1})-p_{j}^{2}(T^{1}_{k^{\prime}})](a-b)\geq 0.\end{split}$ (G-4)
To better analyze whether (G-4) is true, we separate (G-4) into three parts, which are given by
$\begin{split}u_{k}^{1}(T_{k}^{1})-u_{k}^{1}(T^{1}_{k^{\prime}})&+u^{1}_{k^{\prime}}(T^{1}_{k^{\prime}})-u^{1}_{k^{\prime}}(T^{1}_{k})\\ &\quad=(\theta_{k}^{1}-\theta^{1}_{k^{\prime}})(R_{k}^{1}-R^{1}_{k^{\prime}}),\end{split}$ (G-5)
$p_{j}^{2}(T_{k}^{1})-p_{j}^{2}(T^{1}_{k^{\prime}}),$ (G-6)
@@
$a-b=\theta_{j}^{2}c\sum_{l=1}^{j-1}\Bigg[\bigg(\frac{1}{\theta_{l}^{2}}-\frac{1}{\theta_{l+1}^{2}}\bigg)(T_{l}^{2}(\theta_{k}^{1})-T_{l}^{2}(\theta^{1}_{k^{\prime}}))\Bigg].$ (G-7)
Since the types of period 1 and period 2 are positively correlated [7, 14], we have $p_{j}^{2}(T_{k}^{1})>p_{j}^{2}(T^{1}_{k^{\prime}})$, i.e., (G-6) is positive. To make (G-4) hold, (G-5) and (G-7) must be positive. Following Lemma 7, (G-5) will be positive if $T_{k}^{1}>T^{1}_{k^{\prime}}$, and (G-7) will be positive if $T_{j}^{2}(T_{k}^{1})>T_{j}^{2}(T^{1}_{k^{\prime}})$. Thus, the proof of Lemma 7 is completed.
Appendix H Proof for Lemma 8
@@
$u_{k}^{1}(T_{k}^{1})+\beta\sum_{j=1}^{K}p_{j}^{2}(T_{k}^{1})[u_{j}^{2}(T_{j}^{2}(T_{k}^{1}))-u_{j}^{2}(T_{j}^{2}(T^{1}_{k^{\prime}}))]\geq u_{k}^{1}(T^{1}_{k^{\prime}}).$ (H-1)
Since $\beta$ and $p_{j}^{2}(T_{k}^{1})$ are small, (H-1) can be simplified into
$u_{k}^{1}(T_{k}^{1})\geq u_{k}^{1}(T^{1}_{k^{\prime}}),$ (H-2)
@@
$\theta_{k}^{1}(R_{k}^{1}-R^{1}_{k^{\prime}})\geq c(T_{k}^{1}-T^{1}_{k^{\prime}}).$ (H-3)
We first prove the sufficiency of Lemma 8, namely if $T_{k}^{1}>T^{1}_{k^{\prime}}$, then $R_{k}^{1}>R^{1}_{k^{\prime}}$. From (H-3), we can know that $R_{k}^{1}>R^{1}_{k^{\prime}}$ since $T_{k}^{1}>T^{1}_{k^{\prime}}$. Thus, the sufficiency is proved.
@@
$u^{1}_{k^{\prime}}(T^{1}_{k^{\prime}})\geq u^{1}_{k^{\prime}}(T^{1}_{k}),$ (H-4)
@@
$\theta^{1}_{k^{\prime}}(R^{1}_{k^{\prime}}-R_{k}^{1})\geq c(T^{1}_{k^{\prime}}-T^{1}_{k}).$ (H-5)
Given $R^{1}_{k^{\prime}}>R_{k}^{1}$, the left-hand side of (H-5) is negative. Thus, we have $T^{1}_{k}-T^{1}_{k^{\prime}}$. The proof of Lemma 8 is completed.